solve the proportion for x: 3/4/9=8/x please show steps
Answer by Jk22(389)

3/4/9=8/x, the operation go from left to right :
(3/4)/9 = 8/x

3/(4*9) = 8/x, we can simplify the LHS by 3 :

1/12 = 8 / x, and augment the LHS by 8 :

8/96 = 8 / x,

hence x = 96